8. Curved surface area of a cone is 308 cm2 and its slant height is 14 cm. Find
(i) radius of the base and


SURFACE MEAS AND VOLUMES 189
(ii) total surface area of the cone.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i') Given: wl = 308 72'2 x r >< 14 = 308 308 r = -= 44 7 cm .. Radius of the base = 7 cm. (ii) 'lbtal surface area = wl + n:r2 = 308 +722 >< (7'f = 462 cm2